OpenStudy (maddog3354):

Amount Financed (m) = $1,400 Number of Payments per year (y) = 12 Number of Payments (n) = 24 Total Interest (c) = $200.50 Interest = ____%?

OpenStudy (tkhunny):

You MUST show YOUR work.

OpenStudy (tkhunny):

If m is the amount financed and P is the periodic payment, \(m = P\cdot\dfrac{1 - v^{n}}{i}\). This is the standard formula for periodic payment over time for level interest. You are given m = 1400 You are given n = 24 You are ALMOST given P. You can figure it out. P = (1400 + 200.50)/24 = 1600.50/24 = 66.69 You're almost done. There are two things yet to do.